# SAM 2++: Tracking Anything at Any Granularity

### v3

@@ 4.3 Training and Inference Details
Training. SAM 2++ performs multi-task training on tracking tasks with different granularity (mask, box, point), initialized from SAM 2 base. We decoupled the memory-related modules, including creating a separate copy of the memory encoder and implementing dedicated LoRA parameters for memory attention for each task. In addition to our Tracking-Any-Granularity dataset, we adopt DAVIS 2017 [davis17], YoutubeVOS 2019 [youtube-vos] and MOSE [MOSE] for the mask task; LaSOT [lasot], GOT10k [got10k], TrackingNet [TrackingNet] and COCO [coco] for the box task; TapVid Kinetics [tapvid], PointOdyssey [zheng2023pointodyssey], and PerceptionTest [patraucean2023perception] for the point task. During training, we use 8-frame sequences with up to 3 targets in the first frame, The first frame and one randomly selected frame serve as conditional frames, receiving either normal prompts or interactive prompts with equal probability. For further details, please refer to the Appendix.
Losses and optimization. For mask tracking, we combine focal and dice losses for mask prediction, L1 loss for IoU prediction, and cross-entropy loss for occlusion prediction. Box tracking extends it with ciou [zheng2020diouciou] and L1 losses for box predictions. Point tracking adds L1 loss for point predictions with soft argmax. In multi-prediction scenarios, we only supervise the task prediction with the lowest combined loss while supervising all IoU predictions. For occluded frames, denoted as $1-\mathbbm{1}_{obj}$, we skip the supervision of the task results and IoU prediction, but maintain occlusion prediction supervision. In summary, the multi-task training loss can be written as:
$\mathcal{L}_{Mask}$ $=(\lambda_{mask}^{focal}\mathcal{L}_{mask}^{focal}+\lambda_{mask}^{dice}\mathcal{L}_{mask}^{dice})\times\mathbbm{1}_{obj}$ (6)
$+\lambda_{IoU}^{L1}\mathcal{L}_{IoU}^{L1}\times\mathbbm{1}_{obj}+\lambda_{obj}^{CE}\mathcal{L}_{obj}^{CE}$
$\mathcal{L}_{Box}$ $=\mathcal{L}_{Mask}+(\lambda_{box}^{ciou}\mathcal{L}_{box}^{ciou}+\lambda_{box}^{L1}\mathcal{L}_{box}^{L1})\times\mathbbm{1}_{obj}$
$\mathcal{L}_{Point}$ $=\mathcal{L}_{Mask}+\lambda_{point}^{L1}\mathcal{L}_{point}^{L1}\times\mathbbm{1}_{obj}$
Inference. During inference, we follow a fully online inference setting where only the ground truth of the first frame serves as the initial prompt without any subsequent corrections and future information. Our model operates on full frames without post-processing strategies like center cropping, which is commonly used in tracking tasks.
5 Data
Figure 5: Statistics on sources and attributes distribution of Tracking-Any-Granularity Dataset. The link in (c) reflects the frequent co-occurrence of multiple attributes in a sequence.
Table 1: Comparison of our datasets with public datasets of three tracking datasets in terms of videos, duration, and annotations.
[…]
BURST [burst] 2914 1734 (0.60) 624240 (214) $\geq$480p 6 600157 (206) × × Semi-Automatic Multi-Task
LVOS [LVOS_V1] 220 351 (1.60) 126280 (574) 720p 6 156432 (711) × × Manual Long-term
LVOS v2 [LVOS_V2] 720 823 (1.14) 296401 (412) 720p 6 407945 (567) × × Manual Large-scale, long-term
MOSE [MOSE] 2149 443.62 (0.21) $\sim$159600 (73) 1080p 6 431725 (201) × × Semi-Automatic Complext scenarios
YoutubeVOS-19 [youtube-vos] 4453 334.8 (0.08) 120532 (27) 720p 6 197272 (44) × × Manual Large-scale
VOST [vost] 713 252 (0.35) 75547 (106) 1080p 5 175913 (247) × × Semi-Automatic Object transmission
LaSOT [lasot] 1400 1950 (1.39) 3.52M (2506) 720p 30 × 3.52M × Manual Large-scale, long-term
[…]
TrackingNet [TrackingNet] 30643 8400 (0.27) 14431266 (471) 360p 30 × 14431266 × Semi-Automatic Large-scale
UAV123 [uav123] 123 62.5 (0.51) 112578 (915) 720p 30 × 112578 × Semi-Automatic Unmanned aerial vehicles
NfS [NFS] 100 26.58 (0.27) 383K (3830) 720p 240 × 383K × Manual High Frame Rate
OTB-100 [otb] 100 32.8 (0.33) 59040 (590) $\geq$360p 30 × 59040 × Manual Real world
TNL2K [wang2021tnl2k] 2000 691.3 (0.35) 1244340 (622) 720p 30 × 1244340 × Manual Language-based
VastTrack [vasttrack] 50610 11664 (0.23) 4.2M (83) 480p-720p 6 × 4.2M × Manual Abundant categories
[…]
PointOdyssey [zheng2023pointodyssey] 104 120 (1.15) $\sim$216K (2035) 540p 30 × × 49B (0.471B) Automatic Real world, long-term
TAP-Vid Kinetics [tapvid] 1189 198.17 (0.17) 297250 (250) $\geq$720p 25 × × 4725959 (3974) Semi-Automatic Abitrary point
TAP-Vid DAVIS [tapvid] 30 - (-) 1999 (66.6) 1080p - × × 28824 (960.8) Semi-Automatic Abitrary point
TAP-Vid RGB-Stacking [tapvid] 50 - (-) 12500 (250) 256x256 - × × 303436 (6068.7) Semi-Automatic Abitrary point
[…]
We developed a comprehensive dataset for training our unified model, termed Tracking-Any-Granularity (TAG), with annotations across three granularities: segmentation masks, bounding boxes, and key points. Our dataset contains 6,000 high-resolution videos featuring diverse scenes, objects, and challenging scenarios (e.g., occlusion, motion blur, etc.). With a three-phase data engine with model-in-the-loop annotation workflows and strict multi-stage quality checks, we ensure large-scale, high-quality, and consistent annotations. For more details, please refer to the Appendix.
5.1 Annotation Pipeline
We designed a coarse-to-fine annotation pipeline to ensure high-quality multi-granularity annotations as demonstrated in Fig. 3. Firstly, we collected videos from YouTube that meet our quality standards and exhibit diverse tracking challenges. Then comes the coarse annotation stage, where annotators mark key points and tight bounding boxes on target objects. Next, in the fine annotation stage, we leverage SAM to generate initial masks from coarse annotations, which annotators then refine. Experts perform quality checks throughout to ensure annotation consistency and accuracy, particularly for challenging scenarios like occlusions and motion blur. As for the Final Completion stage, the experts check the consistency of the three labellings.
5.2 Data Engine
As shown in Table 2, the Tracking-Any-Granularity dataset is annotated across three phases: 1) Phase ①: Manual annotation of every frame, totaling 1,000 videos. 2) Phase ②: Manual annotation of every 10 frames, totaling 2,000 videos. 3) Phase ③: Manual annotation of every 20 frames, totaling 3,000 videos. In Phases ② and ③, we integrated SAM 2++, which is trained on public datasets and previous phase annotations, to automatically annotate frames between manual-annotated frames. In detail, we divided videos into clips where both first and last frames were manually annotated, then used the annotation of the first frame in each clip as input to infer intermediate frames. To improve annotation quality, we implemented two optional enhancements: (1) performing backward tracking and fusing results with forward tracking, and (2) using the first video frame (guaranteed to contain the target) as an additional starting point when targets might be absent in keyframes. We evaluated these enhancements on validation data in Phase ① to select optimal strategies for each tracking task.
5.3 Tracking-Any-Granularity Dataset
Compared with existing datasets in video tracking tasks, our Tracking-Any-Granularity dataset stands out as the only one providing annotations at all three granularities simultaneously. We compare our dataset with numerous public datasets in Table 1, showing our dataset contains significantly more videos and annotations than they do, creating a substantial resource for multi-granularity tracking research. Fig. 4 shows examples from our dataset, annotated at all three granularities and exhibiting diverse challenges.
Table 2: Evolution of data engine phases, showing the interval and number of manual annotations.
TAG Videos Interval Points Boxes/Masks Total Frames Total Len (s)
Phase ① 1,000 1 523,137 348,715 354,625 12,540.9
Phase ② 2,000 10 87,708 75,923 787,643 28,164.5
Phase ③ 3,000 20 60,809 53,917 1,058,623 39,617.1
Total 6,000 - 671,654 478,555 2,200,891 80,322.4
Table 3: State-of-the-art comparison on Video Object Segmentation Task.
[…]
STCN [stcn] 45.8 45.6 45.9 62.0 58.5 65.4 50.1 46.1 54.1 70.4 65.9 75.0 76.2 72.2 80.2 32.7 22.7 82.8 81.3 85.6 78.3 86.0
AOT-SwinB [aot] 54.9 55.3 54.8 73.8 70.1 77.4 60.2 56.2 64.1 78.1 73.1 83.2 80.9 76.4 85.4 39.3 29.4 85.3 84.6 89.5 79.3 87.7
DeAOT-SwinB [deaot] 58.5 58.0 58.6 72.1 68.5 75.8 61.7 57.6 65.9 79.6 74.8 84.4 81.6 77.3 85.9 43.0 29.4 86.4 85.4 90.3 80.6 89.3
XMem [xmem] 52.3 51.9 52.3 64.7 61.8 67.5 56.3 52.1 60.5 74.4 70.1 78.6 75.7 71.8 79.6 37.9 25.0 85.5 84.3 88.6 80.5 88.7
DEVA [deva] 56.1 55.4 56.2 72.2 68.6 75.9 66.4 62.2 70.6 77.9 73.1 82.6 82.1 78.0 86.1 40.9 27.3 86.3 85.3 89.8 80.6 89.2
Cutie-base+ [cutie] 55.8 55.6 55.8 71.4 68.6 74.3 66.2 62.3 70.1 79.0 75.0 83.0 83.8 80.0 87.7 44.7 32.7 86.9 86.2 90.7 81.6 89.2
[…]
JointFormer [jointformer] - - - 71.7 68.8 74.7 69.7 65.8 73.6 76.6 72.8 80.5 79.1 75.5 82.7 - - 87.0 86.1 90.6 82.0 89.5
[…]
Table 4: State-of-the-art comparison on Single Object Tracking Task.
Methods GOT10ktest $_{SOT}^{test} TAG$$_{SOT}^{val}$ TrackingNet TNL2K VastTrack
AO $SR_{0.5}$ $SR_{0.75}$ AUC $_{Norm} PP$ AUC $_{Norm} PP$ AUC $_{Norm} PP$ AUC $_{Norm} PP$ AUC $_{Norm} PP$
OSTrack [ye2022joint_OStrack] 74.8 84.4 72.7 69.7 78.8 69.9 68.3 77.1 66.2 83.9 88.5 83.2 57.6 74.4 58.9 33.7 40.8 31.4
SimTrack [pauwels2015simtrack] 71.1 80.5 68.1 64.1 72.4 60.5 65.8 73.7 63.7 82.3 - 86.5 54.4 70.2 53.7 34.5 40.5 30.4
[…]
DropTrack [wu2023dropmae_droptrack] 76.8 86.9 74.4 71.1 80.5 72.1 70.8 80.4 69.4 83.8 88.5 83.1 58.5 75.7 60.3 37.5 45.9 36.4
GRM [Gao_2023_CVPR_GRM] 73.1 82.3 71.4 69.1 77.4 69.1 68.5 77.3 66.5 84.0 88.7 83.3 56.8 73.2 57.7 34.6 42.2 32.3
SeqTrack [chen2023seqtrack] 77.0 85.8 76.1 69.8 79.4 71.5 68.5 78.2 67.8 83.9 88.8 83.6 57.8 75.3 60.8 35.8 44.8 35.3
ARTrack [wei2023autoregressiveARtrack] 76.8 85.8 75.7 71.1 78.7 70.9 69.9 76.9 67.1 84.2 88.7 83.5 57.9 73.9 59.6 35.7 42.1 32.4
ARTrack-V2 [bai2024artrackv2] 76.3 85.5 74.3 71.8 79.5 71.9 70.2 78.0 68.3 84.9 89.3 84.5 57.1 73.4 58.7 37.0 44.5 34.8
ROMTrack [cai2023robust_romtrack] 75.6 85.4 73.7 71.3 80.8 72.8 69.2 77.8 68.5 84.1 89.0 83.7 58.2 75.3 59.8 37.1 45.5 36.2
HIPTrack [cai2024hiptrack] 78.2 88.5 76.6 71.4 81.0 72.5 72.0 82.2 71.0 84.5 89.1 83.8 59.8 77.1 62.1 38.6 46.3 36.8
LoRAT [lin2024tracking_lorat] 75.1 84.8 74.4 70.5 79.7 68.7 72.7 82.2 74.4 83.5 87.9 82.1 58.8 76.2 61.4 38.7 41.1 37.8
[…]
Scene and Attribute. To enable a more comprehensive analysis of tracking approaches, it is critically important to identify video scenes and attributes of our dataset. Fig. 5 demonstrates that our dataset encompasses a diverse range of sources, highlighting its robust diversity and enabling it to serve as a powerful benchmark for evaluating tracking performance across various environments. Furthermore, we label each sequence with 18 attributes that represent various video challenges. It is worth noting that these attributes are not mutually exclusive, and a single video may contain multiple challenges. Fig. 5(a) and (b) illustrate the distribution of challenges in each video and their mutual dependencies. Motion Blur, Deformation, and Partial Occlusion are the most common challenges in our dataset, demonstrating its high level of difficulty. We further explore the likelihood of videos being linked to multiple attributes, and Fig. 5(c) indicates that most videos possess more than one attribute.
Dataset Splits. We selected 150 validation videos and 150 test videos from the 1,000 fully annotated videos in Phase ① with stratified sampling based on both category and source, which ensures a balanced distribution.
6 Experiments
6.1 Comparison to state-of-the-art on three tasks
Video Object Segmentation. The comparisons between our model and previous semi-supervised VOS methods are demonstrated in Table 3, including YoutubeVOS-19 [youtube-vos], MOSE [MOSE], LVOS-v2 [LVOS_V2], BURST [burst], VOST [vost], VISOR [visor], and our TAG dataset. We use the standard metric $\mathcal{J\&F}$ [vos-benchmark] that averages Jaccard index and contour accuracy in most benchmarks, but adopt Higher Order Tracking Accuracy (HOTA) [hota] in the BURST benchmark. Results show that our model outperforms individual video object segmentation models.
Single Object Tracking. We compare the performance of our proposed model on three benchmarks in Table 4, including TrackingNet [TrackingNet], GOT-10k [got10k], TNL2K [wang2021tnl2k], VastTrack [vasttrack] and our TAG dataset, and all compared models are trained on four datasets. We choose the Average Overlap (AO) for the GOT-10k benchmark, and Area Under the Curve (AUC) for the other benchmarks. Experiments demonstrate that our model consistently outperforms existing methods across nearly all benchmarks.
Table 5: State-of-the-art comparison on Point Tracking Task. model and model means offline and modified online trackers.
Method Key Point Arbitrary Point
BADJA PerceptionTestval $_{PT}^{test} TAG$$_{PT}^{val}$ Tapviddavis Tapvidrgb
pips [harley2022pips] 64.2 41.5 19.0 19.8 40.9 28.5
pips [harley2022pips] 56.9 27.2 12.3 12.2 26.3 20.8
pips++ [zheng2023pointodyssey] 56.6 62.6 20.9 23.1 59.8 58.5
CoTracker [cotrackerv1] 65.3 59.8 23.3 22.3 60.9 63.1
CoTracker [cotrackerv1] 55.1 48.9 18.8 18.1 50.8 46.1
CoTracker3 [cotrackerv3] 72.7 71.3 29.6 29.1 65.6 70.6
CoTracker3 [cotrackerv3] 66.3 66.3 25.8 24.9 59.2 63.6
TAPTR [TAPTR] 69.1 59.4 23.7 23.8 61.2 58.0
TAPTR [TAPTR] 63.0 48.9 20.4 19.0 52.3 39.4
TAPIR [doersch2023tapir] 64.2 59.6 21.3 24.6 56.8 50.4
LocoTrack [LocoTrack] 68.7 67.1 25.2 30.2 62.7 69.2
Track-On [Track-On] 69.7 69.7 24.8 25.8 64.5 64.5
[…]
Figure 6: Visualization of memory design at different components and granularities. In the visualization of each component, the left side represents our task-adaptive memory mechanism and the right side represents full parameter sharing.
Online Point Tracking. We compare our method to prior works in Table 5 on four benchmarks, including BADJA [badja] and Perception Test [patraucean2023perception] for key point tracking, TAP-Vid [tapvid] for arbitrary point tracking, and our TAG dataset in the ‘query first’ evaluation, which means points appearing in the first frame are used as queries. We report the Percentage of Correct Keypoint-Transfer (PCK-T) for the BADJA benchmark, and Average Jaccard (AJ) for the remaining benchmarks. However, most of the current methods are offline trackers, which process long-temporal window frames or even the entire video to be able to see the future frame, and do not match the online setup required by real applications. For a fair comparison, we modified their input so that there is no future information inside the window, denoted as model, to enable inference online. Experimental results show a substantial decrease in model performance when the input data is switched from offline to online. The comparative analysis reveals the effectiveness of our approach on keypoint tracking benchmarks, which surpasses competing models. Furthermore, although our model is trained on keypoint datasets, it demonstrates generalization capability on arbitrary point tracking datasets.
Table 6: Analysis of Mixed training strategy and Data Engine. ‘Shared’ and ‘Decoupled’ denote that the module shares parameters or decouples parameters between different tasks, respectively
[…]
$\times$ Shared Shared 69.3 84.7 28.1 26.1 64.1 30.3
$\times$ Shared Decoupled 73.6 86.4 64.5 57.3 63.0 28.7
+① Freezed Decoupled 74.2 87.1 67.8 59.2 68.8 29.9
+① Decoupled Decoupled 75.4 86.8 66.9 58.5 72.1 32.7
+① Shared Decoupled 76.4 87.1 68.9 58.2 71.9 33.1
[…]
6.2 Exploration Studies
Study on Mixed training strategy. To verify the effectiveness of task-adaptive memory during multi-task joint training, we compare the results of single-task training with different parameter settings during multi-task mixing training. As shown in Table 6, when a single set of parameters is naively shared for multi-task joint training, the differences between tasks lead to a performance drop across all tasks. This indicates that the encoding and retrieval components of the memory module need to be decoupled for different tasks. In addition, when the image encoder is either frozen or similarly decoupled, the performance is inferior to a shared encoder. This suggests that the image encoder benefits from exposure to more data and is not adversely affected by task differences.
Study on data engine. To validate the effectiveness of our data engine, we evaluated the performance when trained on different phases of our TAG dataset, as shown in Table 6. The results demonstrate that our proposed dataset enhances the performance on other datasets, indicating high diversity and generalizability. After training with more data from phases ② and ③, the performance is further improved across all three tasks, demonstrating the effectiveness of the supplementary data provided by our data engine.
6.3 Visualization
7 Conclusion
We present SAM 2++, a foundational model for tracking targets at any granularity, built upon three key contributions: 1) Unifying task processing through task-specific prompts for inputs and a Unified Decoder for outputs; 2) Unifying task states across different granularities via a task-adaptive memory mechanism; 3) Introducing the Tracking-Any-Granularity dataset for training and benchmarking video tracking at multiple granularities. We hope that SAM 2++ can serve as a strong baseline for general tracking and provide a powerful impetus for future research.
[…]
Supplementary Material
In the appendix, we present a more detailed discussion of the topics covered in the main text, with the specifics of each section described as follows:
  • Section 8: Model Details
  • Section 9: Data Details
  • Section 10: Additional Experiments
  • Section 11: Limitations, Impacts, and Future
8 Model Details
8.1 Model Architecture
Task-Specific Prompt. In order to unify the different inputs for each task and not modify the structure of the original Prompt Encoder, we provide task-specific prompt for each task, which provides an accurate and efficient representation of the target state of each task. The design of the task-specific prompt for each task is as follows:
  • Mask tracking: {0, 1} mask to accurately describe the shape and boundaries of the target;
  • Box tracking: bounding boxes in the top-left and bottom-right corners;
  • Point tracking: Besides the exact point coordinates, we provide a (0, 1) Gaussian mask generated from the points to better represent the target in memory and align with the mask outputs from the Decoder.
Unified Decoder. We made minor modifications to the Mask Decoder to obtain the desired outputs for each task. Specifically: (1) we added a Corner Head for the box tracking task to directly output the bounding box, thereby avoiding the low precision and lack of gradients associated with the outer box operation; and (2) we applied an argmax operation to the masked output (or a soft-argmax operation during training to ensure gradient flow) to obtain the point coordinates, which are aligned with the Gaussian mask of the prompts.
Task-adaptive Memory. Based on the analysis described in the main text, we decouple the memory components for each task. Specifically, for Memory Attention in the Transformer architecture, we configure an independent LoRA for each task. For the Memory Encoder in the convolutional architecture, we set up a separate copy for each task, resulting in only a minimal increase in parameters.
8.2 Training and Inference Details
Table 7: Hyperparameters and details of SAM 2++ training in three tasks.
@@ 8.2 Training and Inference Details
LaSOT,
COCO,
GOT-10K,
TrackingNet,
@@
TAP-Vid Kinetics,
PerceptionTest,
PointOdyssey ,
@@
batch size 16
drop path 0.2 (B+)
epochs 150
resolution 1024
precision bfloat16
optimizer AdamW
optimizer momentum $\beta_{1}$, $\beta_{2}$ =0.9, 0.999
gradient clipping type: L2 , max: 0.1
weight decay 0.1
learning rate (lr) backbone: 5.0e-6, other: 3.0e-06
lr schedule cosine
layer-wise decay 0.9 (B+)
@@
hflip,
affine (deg: 25, shear: 20) in mask, w/o affine in box and point,
resize to 1024 (square),
colorjitter (b: 0.1, c: 0.03, s: 0.03, h: null),
grayscale (0.05),
per frame colorjitter (b: 0.1, c: 0.05, s: 0.05, h: null)
mask losses focal (20), dice (1) focal (0.5), dice (0.1) focal (20), dice (1)
@@
Iou loss l1 loss (1)
occlusion loss cross-entropy (1)
input prompt
mask (0.5),
noisy box (0.25),
sampled points (0.25)
box (0.5),
noisy box (0.25),
sampled points (0.25)
point coordinate & Gaussian mask (1)
@@
Task-specific Modification
Decoupled & LoRA Memory Attention, Decoupled Memory Encoder,
Shared & LoRA Image Encoder
Other Setting - Corner Head
Gaussian mask settings
[…]
Training. The SAM 2++ training is conducted on 16 H800 GPUs. We expand SAM 2++ to three tasks: semi-supervised video object segmentation (mask), single object tracking (box), and online point tracking (point), by processing input, prompt, memory, and output into a unified format used by SAM 2. Our training process is based on the mask tracking task in SAM 2 and we make minimal modifications to it while adding task-specific requirements from the other tasks. Table 7 describes the training settings for three tasks in detail, and other settings not mentioned follow SAM 2.
Training is performed jointly on data of the three tasks. In addition to our Tracking-Any-Granularity dataset, we used DAVIS-17 [davis17], YoutubeVOS-19 [youtube-vos] and MOSE [MOSE] for the mask task, LaSOT [lasot], GOT-10k [got10k], TrackingNet [TrackingNet] and COCO [coco] for the box task, and TAP-Vid Kinetics [tapvid], PointOdyssey [zheng2023pointodyssey], and PerceptionTest [patraucean2023perception] for the point task. To enable the model to be simultaneously capable of all three tasks and to optimize training efficiency, we adopt the strategy of alternating between the three tasks. Specifically, we implement parallelisation by sampling a whole batch at each step of training, which is entirely derived from the data of a particular task. The sampling probability is set to 1:4:5 to balance the performance of the three tasks.
@@
In addition, if interactive input is used as initial prompt, we select up to 2 frames as corrective frames to add corrective clicks on them: after predicting the selected frame, we sample a positive point from the false positive region between the prediction and ground truth or a negative point from the false negative region as a corrective point, and use it as additional prompt to get a new prediction along with all previous cumulative prompt from that frame. This operation is repeated until 7 corrective points have been added. In addition, if the box tracking task uses the box format to compute the regional differences between the prediction and the GT, there is an overwhelming problem that the sampled corrective clicks may fall at the boundaries of the box instead of inside the target, which is contrary to the actual interaction. Therefore, we choose to compute the difference in mask format, and use SAM 2 and sam-hq [sam_hq] with box annotations to obtain the pseudo-GT mask on SOT datasets because of its good segmentation ability.
Losses and optimization. Following the mask tracking task in SAM 2, we adopt the linear combination of focal loss $\mathcal{L}_{mask}^{focal}$ and dice loss $\mathcal{L}_{mask}^{dice}$ for the mask prediction, L1 loss for the IoU prediction $\mathcal{L}_{L1}^{IoU}$, and cross-entropy loss for object occlusion prediction $\mathcal{L}_{CE}^{obj}$. During the box tracking task, we adopt the corner head to predict the bounding boxes and add additional ciou loss [zheng2020diouciou] and L1 loss to supervise the box prediction. As for the point tracking task, we select the highest probability position from the mask prediction as point prediction, and use soft argmax [softargmax] during training for making the process derivable instead of the undifferentiable argmax function. Beyond the loss on mask in the form of Gaussian map, we add an L1 loss between the prediction and ground-truth point to directly optimize the distance and accuracy of the points. For multi-prediction scenario, we only supervise the task predictions (masks, boxes, and points) with the lowest loss, which is a combination of $\mathcal{L}^{mask}$, $\mathcal{L}^{box}$ and $\mathcal{L}^{point}$, but supervise the IoU predictions of all task predictions to learn to synchronise the quality of predictions. Furthermore, if the target is missing in some frames due to disappearance or cropping, we do not supervise the task predictions or iou predictions on them in all three tasks, but always supervise the occlusion prediction from an MLP head, no matter if the ground-truth exists or not. In summary, the supervision losses for the three tasks can be written as:
@@
where $\mathbbm{1}_{obj}$ denotes we supervise task and IoU prediction only if the object exists, and $\lambda$ represents the weights of different losses. The specific hyperparameters for the training are shown in Table 7.
@@
8.3 Efficiency Analysis
Table 8: Hyperparameters and details of SAM 2++ training in three tasks.
Modules GFlops Main Param Lora Param
Image Encoder 264.4 69.1 22.25
Memory Encoder 5.0 1.4 $\times$ 3 -
Mask Decoder 53.4 9.9 -
Memory Attention 27.4 5.9 4.3 $\times$ 3
Total 350.2 89.1 35.1
To present a comprehensive view of the model’s computational complexity and parameter overhead, we provide a detailed breakdown of the computational cost for each module in Table 8, including GFLOPs, the number of parameters, and the LoRA parameters introduced during training. We would like to kindly note that the LoRA parameters exist only during training and are merged into the main model weights at inference time. Therefore, they do not introduce additional parameters or computational overhead during inference. In addition, although the model contains multiple Memory Encoders designed for different granularities, only the branch corresponding to the current granularity is activated at inference, while the others remain inactive (excluded from computation), ensuring that no extra inference cost is incurred.
9 Data Details

### v4

@@ 4.3 Training and Inference Details
Training. The SAM 2++ training is conducted on 16 H800 GPUs. We expand SAM 2++ to three tasks: semi-supervised video object segmentation (mask), single object tracking (box), and online point tracking (point), by processing input, prompt, memory, and output into a unified format used by SAM 2. Our training process is based on the mask tracking task in SAM 2 and we make minimal modifications to it while adding task-specific requirements from the other tasks. Table I describes the training settings for three tasks in detail, and other settings not mentioned follow SAM 2. We initialize SAM 2++ from SAM 2 base and decouple the memory-related modules, including creating a separate copy of the memory encoder and implementing dedicated LoRA parameters for memory attention for each task.
TABLE I: Hyperparameters and details of SAM 2++ training in three tasks.
@@
DAVIS 2017, MOSE, YoutubeVOS 2019
@@
LaSOT, COCO, GOT-10K, TrackingNet
@@
shared optimization settings
batch size: 16; drop path: 0.2 (B+); epochs: 150; resolution: 1024; precision: bfloat16; optimizer: AdamW; $\beta_{1}$, $\beta_{2}$: 0.9, 0.999;
gradient clipping: L2, max 0.1; weight decay: 0.1; lr: backbone 5.0e-6, other 3.0e-06; schedule: cosine; layer-wise decay: 0.9 (B+)
@@
hflip; affine (deg: 25, shear: 20) in mask, w/o affine in box and point; resize to 1024 (square);
colorjitter (b: 0.1, c: 0.03, s: 0.03, h: null); grayscale (0.05); per-frame colorjitter (b: 0.1, c: 0.05, s: 0.05, h: null)
Loss mask losses focal (20), dice (1) focal (0.5), dice (0.1) focal (20), dice (1)
@@
Iou loss l1 loss (1) l1 loss (1) l1 loss (1)
occlusion loss cross-entropy (1) cross-entropy (1) cross-entropy (1)
input prompt mask (0.5), noisy box (0.25), sampled points (0.25) box (0.5), noisy box (0.25), sampled points (0.25) point coordinate & Gaussian mask (1)
@@
Task-specific Modification Decoupled & LoRA Memory Attention; Decoupled Memory Encoder; Shared & LoRA Image Encoder
Other Setting - Corner Head $\left\{\begin{array}[]{l}\text{Ep }0\sim 20:\ \text{radius}=50,\ \text{sigma}=16\\ \text{Ep }20\sim 50:\ \text{radius}=20,\ \text{sigma}=8\\ \text{Ep }50\sim 100:\ \text{radius}=5,\ \text{sigma}=2\end{array}\right$.
Training is performed jointly on data of the three tasks. In addition to our Tracking-Any-Granularity dataset, we used DAVIS-17 [56], YoutubeVOS-19 [73] and MOSE [21] for the mask task, LaSOT [25], GOT-10k [34], TrackingNet [51] and COCO [47] for the box task, and TAP-Vid Kinetics [23], PointOdyssey [86], and PerceptionTest [52] for the point task. To enable the model to be simultaneously capable of all three tasks and to optimize training efficiency, we adopt the strategy of alternating between the three tasks. Specifically, we implement parallelisation by sampling a whole batch at each step of training, which is entirely derived from the data of a particular task. The sampling probability is set to 1:4:5 to balance the performance of the three tasks.
@@
In addition, if interactive input is used as initial prompt, we select up to 2 frames as corrective frames to add corrective clicks on them: after predicting the selected frame, we sample a positive point from the false positive region between the prediction and ground truth or a negative point from the false negative region as a corrective point, and use it as additional prompt to get a new prediction along with all previous cumulative prompt from that frame. This operation is repeated until 7 corrective points have been added. In addition, if the box tracking task uses the box format to compute the regional differences between the prediction and the GT, there is an overwhelming problem that the sampled corrective clicks may fall at the boundaries of the box instead of inside the target, which is contrary to the actual interaction. Therefore, we choose to compute the difference in mask format, and use SAM 2 and sam-hq [39] with box annotations to obtain the pseudo-GT mask on SOT datasets because of its good segmentation ability.
Losses and optimization. Following the mask tracking task in SAM 2, we adopt the linear combination of focal loss $\mathcal{L}_{mask}^{focal}$ and dice loss $\mathcal{L}_{mask}^{dice}$ for the mask prediction, L1 loss for the IoU prediction $\mathcal{L}_{L1}^{IoU}$, and cross-entropy loss for object occlusion prediction $\mathcal{L}_{CE}^{obj}$. During the box tracking task, we adopt the corner head to predict the bounding boxes and add additional ciou loss [87] and L1 loss to supervise the box prediction. As for the point tracking task, we select the highest probability position from the mask prediction as point prediction, and use soft argmax [27] during training for making the process derivable instead of the undifferentiable argmax function. Beyond the loss on mask in the form of Gaussian map, we add an L1 loss between the prediction and ground-truth point to directly optimize the distance and accuracy of the points. For multi-prediction scenario, we only supervise the task predictions (masks, boxes, and points) with the lowest loss, which is a combination of $\mathcal{L}^{mask}$, $\mathcal{L}^{box}$ and $\mathcal{L}^{point}$, but supervise the IoU predictions of all task predictions to learn to synchronise the quality of predictions. Furthermore, if the target is missing in some frames due to disappearance or cropping, we do not supervise the task predictions or iou predictions on them in all three tasks, but always supervise the occlusion prediction from an MLP head, no matter if the ground-truth exists or not. In summary, the supervision losses for the three tasks can be written as
@@
where $\mathbbm{1}_{obj}$ denotes we supervise task and IoU prediction only if the object exists, and $\lambda$ represents the weights of different losses. Table I shows the detailed training settings.
@@
Figure 3: Examples of the Tracking-Any-Granularity dataset. Each sequence is annotated at three granularities: bounding boxes for object-level localization, segmentation masks for pixel-level shape, and key points for fine-grained target tracking.
TABLE II: The interval and number of manual annotations in various phases.
TAG Videos Interval Points Boxes/Masks Total Frames Total Len (s)
Phase ① 1,000 1 523,137 348,715 354,625 12,540.9
Phase ② 2,000 10 87,708 75,923 787,643 28,164.5
Phase ③ 3,000 20 60,809 53,917 1,058,623 39,617.1
Total 6,000 - 671,654 478,555 2,200,891 80,322.4
5 Tracking-Any-Granularity Dataset